OpenStudy (anonymous):

-2x² - 7 = 4x Where are a, b, and c in this equation? How do I solve this with the quadratic equation?

HanAkoSolo (jamierox4ev3r):

remember, a standard quadratic equation is always in this form: \(a^{2}x+bx+c\) so you're going to need to manipulate this equation to set it to this formula

OpenStudy (anonymous):

Is a 2? Is b 4? Is C 7?

OpenStudy (butterflydreamer):

we have this equation: -2x² - 7 = 4x We need to rearrange it so that the x² part is positive. Therefore take the "-2x² - 7" to the other side of the equals sign and we get: -2x² - 7 = 4x 0 = 2x² + 4x +7 Standard form is : ax² + bx + c = 0 Therefore match it up with your equation and we get: a = 2, b = 4 , c = 7 Which you said.. so you're correct ^_^!

OpenStudy (anonymous):

How do I solve this?

OpenStudy (anonymous):

use this formula\[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (butterflydreamer):

yeah use the quadratic formula (as stated by surjithayer) to solve for x :)

OpenStudy (anonymous):

1.58?

OpenStudy (butterflydreamer):

not quite we know a = 2, b = 4 and c = 7 plug these values into the quadratic formula: